# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v6

@@ 2.2 Kolmogorov-Arnold Network generator
Having sampled the intermediary latent variable with ITS such that $z_{q,p}=F^{-1}(\pi_{q,p})(u_{p})$, the generator can be realized flexibly as $\mathcal{G}:\mathcal{Z}\to\mathcal{X}$:
$\bm{\tilde{x}}^{(s)}$ $=\mathcal{G}\left(\bm{z}^{(s)}\right)+\bm{\epsilon}^{(s)},\quad\text{where}\quad z_{q,p}^{(s)}\sim\pi_{q,p}(z),\quad\bm{\epsilon}^{(s)}\sim\mathcal{N}(\bm{\epsilon};\;\bm{0},\sigma_{\epsilon}^{2}\bm{I}),$ (8)
@@
The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}\to\mathbb{R}$.
@@ 2.3 Scaling KAEM
From this point forward, we formulate our derivations over a multivariate probability space $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),P)$, where $\mathcal{Z}$ no longer represents a subinterval of the real line, but rather a subspace of $\mathbb{R}^{(2n_{z}+1)\times n_{z}}$. The measure $P$ defines a product distribution over the latent variables:
$P(\bm{z}\mid\bm{f})=\pi_{1,1}\otimes\cdots\otimes\pi_{2n_{z}+1,n_{z}},$
where each component $\pi_{j,k}$ is defined via exponential tilting, as introduced in Sec. 2.1. The full latent vector $\bm{z}$ is a matrix-valued random variable composed of the independent latent features.
@@ 2.3.1 Generator
The generator, $\mathcal{G}$, can be considered a separate generator network allowing for more flexible realizations, such as convolutional layers when adherence to Eq. 1 is not required. Deepening this generator, as guided by Liu et al. (2024) for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of KAEM:
@@
Adopting a KAN-based generator also allows for training with L1 regularization, followed by pruning with the method proposed by Liu et al. (2024). This enables further model compression beyond the baseline structure defined by KART.
@@ 2.3.2 Mixture prior
As noted in Sec. 2.1, the independence of the priors $\pi_{q,p}$ limits expressivity. Moreover, the summation over $p$ in Eq. 9 accumulates the variances of sampled $z_{q,p}$, which can lead to overly diffuse representations. To address this, particularly for more complex generative tasks such as those in Sec. 3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution, similar to the approach described by Rivera (2024). This involves moving the summation over $p$ out of the generator and absorbing it into the prior, yielding a mixture model:
$\sum_{p=1}^{n_{z}}\pi_{q,p}(z)$ $=\sum_{p=1}^{n_{z}}\alpha_{q,p}\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),\quad\text{where}\quad\sum_{p=1}^{n_{z}}\alpha_{q,p}=1,\quad\alpha_{q,p}\geq 0.$ (13)
Here, $\alpha_{q,p}$ represent the mixture proportions, which must satisfy non-negativity and unit-sum constraints. In this study, they were fixed uniformly as $1/n_{z}$, guided by Rivera (2024). They could alternatively be learned by passing trainable parameters through a softmax function, (to enforce the constraints), however this approach was not used for the experiments in Sec. 3.
The resulting distribution defines a mixture of energy-based priors. Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (Devroye, 2006). However, this violates adherence to KART, since $F$ now corresponds to a discrete rather than continuous CDF.
The mixture proportions are learned by replacing the log-prior in Eq. 19 with the logarithm of Eq.13. To encourage balanced weighting and reduce model complexity, L1 regularization can be applied to $\alpha_{q,p}$ and small values can be pruned.
Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$. Let $u_{\text{mix}}\sim\mathcal{U}(u_{\text{mix}};\;0,1)$. Guided by Devroye (2006) regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
@@
2.3.3 Deep KAN prior
If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach instead of the mixture approach in Sec. 2.3.2. In fact, each $f_{q,p}$ can itself be a realization of Eq. 1 when its domain is restricted to $[0,1]$. This would introduce new challenges:
  1. 1.
    Prior interpretability is reduced, complicating the recovery of trained priors.
  2. 2.
    The component-wise structure is compromised, undermining the validity of ITS and KART.
Instead, we propose removing the summation over $p$ in Eq. 9 and reparameterizing the prior using a multilayer KAN energy function, which can be deepened as needed:
$P\left(\bm{z}\mid\bm{f}\right)=\frac{1}{\bm{Z}}\cdot\exp\left(\sum_{q=1}^{2n_{z}+1}\Psi_{q,p}\left(\sum_{p=1}^{n_{z}}f_{q,p}(z)\right)\right)\cdot\pi_{0}(\bm{z}),$ (15)
where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{n_{z}}$, and is constrained to $0,1]^{n_{z}}$ when strict adherence to KART is required. Sampling from this prior can be carried out using the Unadjusted Langevin Algorithm (ULA), as described in Sec. 2.5.2. This is not investigated in the experiments, since it is simply equivalent to the model of Pang et al. (2020), reformulated as a KAN.
@@ 2.4 Maximum likelihood training
$\min_{\bm{f},\bm{\Phi}}\mathcal{L}(\bm{x}^{(1)},\dots,\bm{x}^{(N_{x})})=\min_{\bm{f},\bm{\Phi}}\left[-\sum_{b=1}^{N_{x}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\mathbf{\Phi})\right)\right].$ (16)
The conditional dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of KAEM. For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq. 16 can be separated out into two log-distributions by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent distribution. This is shown under Sec. A.4.2, and the resulting expectation exhibits a lack of dependence on sampling, as proven under Sec. A.4.3:
$\nabla_{\bm{f},\bm{\Phi}}\left[\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right]$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]+\nabla_{\bm{\Phi}}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\Bigg],$ (17)
@@ 2.4.1 Log-prior
$\log\Big(P\left(\bm{z}\mid\bm{f}\right)\Big)$ $=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (18)
@@
$\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]\Bigg]=\;$ $\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$
-\; $\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (19)
Here, $\bm{z}$ is sampled from $P(\bm{z}\mid\bm{f})$ and $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$ and reflects a univariate element of the prior model. The expectation with respect to the posterior can be approximated using Importance Sampling (IS), outlined under Secs. 2.4.4 &  A.6.1, or a standard Monte Carlo estimator when using Langevin sampling methods. The prior expectation may always be approximated with a Monte Carlo estimator taken across samples:
$\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\quad\text{where}\quad z_{q,p}^{(s)}\sim\pi_{q,p}(z).$ (20)
@@ 2.4.2 Log-likelihood
Depending on the task the log-likelihood can be modeled in multiple ways. For example, a Gaussian model might be adopted for images, and an autoregressive model for text. It can also be realized into a multimodal generative modeling framework using the principles outlined by Yuan et al. (2024).
$\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)$ $=\begin{cases}-\frac{\left\lVert\bm{x}^{(b)}-\bm{\tilde{x}}\right\rVert^{2}_{2}}{2\sigma_{\text{llhood}}^{2}}+\text{const},&\text{(static continuous data)}\\[5.0pt] \sum_{o=1}^{n_{x}}\sum_{c=1}^{K}x_{o,c}^{(b)}\log\tilde{x}_{o,c}+\text{const},&\text{(static discrete data)}\\[5.0pt] \sum_{t=1}^{T}\log\left(P(\bm{x}^{(b,t)}\mid\bm{x}^{(b,1)},\dots,\bm{x}^{(b,t-1)},\bm{z},\bm{\Phi})\right),&\text{(sequential data)}\end{cases}$ (21)
@@
The code in Sec. A.1 also provides an option to augment the first likelihood model with an additional perceptual loss term (Johnson et al., 2016) using features extracted from the pretrained VGG-16 CNN (Simonyan & Zisserman, 2015). This could have yielded higher-fidelity images for Sec. 3.2, but it was not adopted in this study due to the computational overhead of Langevin sampling.
@@ 2.4.3 Prior inverse transform sampling
In Eqs. 9 &  20, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986). ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of KAEM. Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
@@
    CDF inversion: Sampling occurs using the inverse cumulative distribution function (CDF): The CDF for a single component, $F_{\pi_{q,p}}$, is:
@@
    These two integrals can be conveniently and efficiently approximated using Gauss-Kronrod quadrature, (Laurie, 1997), given that $f_{q,p}$ is univariate and defined on a bounded grid:
@@
    where $z_{i}^{\text{node}}$ and $w_{i}$ are the Gauss-Legendre quadrature nodes and weights for the interval $z_{\text{min}},z^{\prime}]$. A random variable is uniformly drawn and shared across $q$, $u_{p}\sim\mathcal{U}(u_{p};\;0,1)$. The first interval where $F_{\pi_{q,p}}(z)$ is greater than $u_{p}$ scaled by $Z_{q,p}$ is identified:
@@
The procedure ensures an efficient, exact and unbiased generation of samples, $z_{q,p}$, distributed according to $\pi_{q,p}$. It can be vectorized to generate samples from all prior elements.
@@ 2.4.4 Posterior Expectations with Importance Sampling
The log-marginal likelihood and its learning gradient presented in Eqs. 16 &  17, involve expectations with respect to the posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is potentially multimodal and complex. We propose Importance Sampling (IS) (Goertzel, 1949) as an efficient method for estimating its expectation when prior task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
@@
    Proposal distribution: $N_{s}$ independent samples, $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}$, are drawn from the proposal distribution, $Q(\bm{z})=P(\bm{z}\mid\bm{f})$, which the target distribution, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$ must be absolutely continuous with respect to. Sampling from $Q$ is tractable with the procedure outlined in Sec. 2.4.3.
@@
    $w_{\text{norm}}(\bm{z}^{(s)})=\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi}))\Big).$ (26)
@@
    Resample: If required, the population, $\{\bm{z}^{(s)},w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N_{s}}$, is resampled from. There are multiple ways to accomplish this, (Douc et al., 2005). Systematic, stratified, and residual resampling are provided in the codebase under Sec. A.1, parallelized on the CPU. In this study, we adopted residual resampling, (explained under Sec. A.6.2).
@@
    Expectation estimation: The posterior expectation of an arbitrary function $\rho(\bm{z})$, (e.g., the log-likelihood), is estimated with a Monte Carlo estimator:
    $\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]\approx{\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})\cdot w_{\text{norm}}(\bm{\bar{z}_{\text{resampled}}}^{(s)})}.$ (28)
This is used to estimate the log-marginal likelihood and its gradient in Eqs. 16 and 17. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.

### v7

@@ 2.2 Kolmogorov-Arnold Network generator
Having sampled the intermediary latent variable with ITS such that $z_{q,p}=F^{-1}_{\pi_{q,p}}(u_{p})$, the generator can be realized flexibly as $\mathcal{G}:\mathcal{Z}\to\mathcal{X}$:
$\bm{\tilde{x}}^{(s)}$ $=\mathcal{G}\left(\bm{z}^{(s)}\right)+\bm{\epsilon}^{(s)},\quad\text{where}\quad z_{q,p}^{(s)}\sim p_{q,p}(z),\quad\bm{\epsilon}^{(s)}\sim\mathcal{N}(\bm{0},\sigma_{\epsilon}^{2}\bm{I}),$ (8)
@@
The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}_{q,p}\to\mathbb{R}$.
@@ 2.3.1 Generator
The generator, $\mathcal{G}$, can be considered a separate generator network allowing for more flexible realizations, such as convolutional layers when adherence to Eq. 1 is not required. Deepening this generator, as guided by liu2024kankolmogorovarnoldnetworks for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of KAEM:
@@
Adopting a KAN-based generator also allows for training with L1 regularization, followed by pruning with the method proposed by liu2024kankolmogorovarnoldnetworks. This enables further model compression beyond the baseline structure defined by KART.
@@ 2.3.2 Mixture prior
As noted in Sec. 2.1, the independence of the priors $\pi_{q,p}$ limits expressivity. Moreover, the summation over $p$ in Eq. 9 accumulates the variances of sampled $z_{q,p}$, which can lead to overly diffuse representations. To address this, particularly for more complex generative tasks such as those in Sec. 3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution, similar to the approach described by rivera2024trainvae. This involves moving the summation over p{p} from 1 to $n_{z}$ out of the generator and absorbing it into the prior, yielding a mixture density:
$p_{\bm{f},q}(z)$ $=\sum_{p=1}^{n_{z}}\alpha_{q,p}\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),\quad\text{where}\quad\sum_{p=1}^{n_{z}}\alpha_{q,p}=1,\quad\alpha_{q,p}\geq 0.$ (13)
Here, $\alpha_{q,p}$ represent the mixture proportions, which must satisfy non-negativity and unit-sum constraints. In this study, they were fixed uniformly as $1/n_{z}$, guided by rivera2024trainvae. They could alternatively be learned by passing trainable parameters through a softmax function, (to enforce the constraints), however this approach was not used for the experiments in Sec. 3.
The result is a mixture of energy-based priors. Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (DEVROYE200683). However, this violates KART, since $F$ now corresponds to a discrete rather than continuous CDF.
The mixture proportions are learned by replacing the log-prior in Eq. 18 with the logarithm of Eq.13. To encourage balanced weighting and reduce model complexity, L1 regularization can be applied to $\alpha_{q,p}$ and small values can be pruned. Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$. Let $u_{\text{mix}}\sim\text{Unif}(0,1)$. Guided by DEVROYE200683 regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
@@ 2.4 Maximum likelihood training
$\min_{\bm{f},\bm{\Phi}}\mathcal{L}(\bm{x}^{(1)},\dots,\bm{x}^{(N_{x})})=\min_{\bm{f},\bm{\Phi}}\left[-\sum_{b=1}^{N_{x}}\log\left(p_{\bm{f},\mathbf{\Phi}}(\bm{x}^{(b)})\right)\right].$ (15)
The dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of KAEM. For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq. 15 can be separated out into two log-densities by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent space. This is shown under Sec. A.4.2, and the resulting expectation exhibits a lack of dependence on sampling, as proven under Sec. A.4.3:
$\nabla_{\bm{f},\bm{\Phi}}\left[\log p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right]$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]+\nabla_{\bm{\Phi}}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\Bigg],$ (16)
@@ 2.4.1 Log-prior
$\log\Big(p_{\bm{f}}\left(\bm{z}\right)\Big)$ $=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (17)
@@
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]\Bigg]=\;$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]-\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (18)
Here, $\bm{z}$ is sampled from $p_{\bm{f}}(\bm{z})$ and $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$ and reflects a univariate element of the prior model. The expectation with respect to the posterior can be approximated using Importance Sampling (IS), outlined under Secs. 2.4.4 &  A.5.1, or a standard Monte Carlo estimator when using Langevin sampling methods. The prior expectation may always be approximated with a Monte Carlo estimator taken across samples:
$\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\quad\text{where}\quad z_{q,p}^{(s)}\sim p_{q,p}(z).$ (19)
@@ 2.4.2 Log-likelihood
Depending on the task the log-likelihood can be modeled in multiple ways. For example, a Gaussian model might be adopted for images, and an autoregressive model for text. It can also be realized into a multimodal generative modeling framework using the principles outlined by yuan2024learningmultimodallatentgenerative.
$\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)$ $=\begin{cases}-\frac{\left\lVert\bm{x}^{(b)}-\bm{\tilde{x}}\right\rVert^{2}_{2}}{2\sigma_{\text{llhood}}^{2}}+\text{const},&\text{(static continuous data)}\\[5.0pt] \sum_{o=1}^{n_{x}}\sum_{c=1}^{K}x_{o,c}^{(b)}\log\tilde{x}_{o,c}+\text{const},&\text{(static discrete data)}\\[5.0pt] \sum_{t=1}^{T}\log\left(p_{\bm{\Phi}}(\bm{x}^{(b,t)}\mid\bm{x}^{(b,1)},\dots,\bm{x}^{(b,t-1)},\bm{z})\right),&\text{(sequential data)}\end{cases}$ (20)
@@
The code in Sec. A.1 also provides an option to augment the first likelihood model with an additional perceptual loss term (johnson2016perceptuallossesrealtimestyle) using features extracted from the pretrained VGG-16 CNN (simonyan2015deepconvolutionalnetworkslargescale). This could have yielded higher-fidelity images for Sec. 3.2, but it was not adopted in this study due to the computational overhead of Langevin sampling.
@@ 2.4.3 Prior inverse transform sampling
In Eqs. 9 &  19, $z_{q,p}$ is drawn from $p_{q,p}$ with inverse transform sampling, (ITS) (devroye1986nonuniform). ITS can be used to sample from a component of the latent density, $p_{q,p}$, made possible given the unique characteristics of KAEM. Specifically, the univariate, finite-grid structure of the KAN-based energy functions enables quadrature/interpolating integral approximations.
@@
    CDF inversion: Sampling occurs using the inverse cumulative density function (CDF): The CDF for a single component, $F_{\pi_{q,p}}$, is:
@@
    These two integrals can be conveniently and efficiently approximated using Gauss-Kronrod quadrature, (gauss-quad), given that $f_{q,p}$ is univariate and defined on a bounded grid:
@@
    where $z_{i}^{\text{node}}$ and $w_{i}$ are the Gauss-Legendre quadrature nodes and weights for the interval $z_{\text{min}},z^{\prime}]$. A random variable is uniformly drawn and shared across $q$, $u_{p}\sim\text{Unif}(0,1)$. The first interval where $F_{\pi_{q,p}}(z)$ is greater than $u_{p}$ scaled by $Z_{q,p}$ is identified:
@@
The procedure ensures an efficient, exact and unbiased generation of samples, $z_{q,p}$, distributed according to $p_{q,p}$. It can be vectorized to generate samples from all prior elements.
@@ 2.4.4 Posterior Expectations with Importance Sampling
The log-marginal likelihood and its learning gradient presented in Eqs. 15 &  16, involve expectations with respect to the posterior $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$, which is potentially multimodal and complex. We propose Importance Sampling (IS) (goertzel1949quota) as an efficient method for estimating its expectation when datasets are simple or when task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
@@
    Proposal distribution: $N_{s}$ independent samples, $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}$, are drawn from the proposal, $q(\bm{z})=p_{\bm{f}}(\bm{z})$, which the target, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$ must be absolutely continuous with respect to. Sampling from the prior is tractable with the procedure outlined in Sec. 2.4.3.
@@
    $w_{\text{norm}}(\bm{z}^{(s)})=\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))\Big).$ (25)
@@
    Resample: If required, the population, $\{\bm{z}^{(s)},w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N_{s}}$, is resampled from. There are multiple ways to accomplish this, (douc2005comparisonresamplingschemesparticle). Systematic, stratified, and residual resampling are provided in the codebase under Sec. A.1, parallelized across samples. In this study, we adopted residual resampling, (explained under Sec. A.5.2).
@@
    Expectation estimation: After resampling, IS weights are uniformly distributed and the posterior expectation of an arbitrary function or density, $\rho(\bm{z})$, (e.g., the log-likelihood), is estimated with a Monte Carlo estimator:
    $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]$ $\approx\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)})\cdot w_{\text{norm}}(\bm{\bm{z}^{(s)})}\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})$. (27)
This is used to estimate the log-marginal likelihood and its gradient in Eqs. 15 and 16. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.